Find the general solution to the given partial differential equation and use it to find the solution satisfying the given initial data.

"\frac{\partial u}{\partial x}=-(2x+y)\frac{\partial u}{\partial y}"

"u(0,y)=1+y^2"


1
Expert's answer
2021-11-03T10:39:03-0400

"\frac{dx}{1}=\frac{dy}{2x+y}"


"(2x+y)dx-dy=0"


"\mu(x)=e^{-x}"


"e^{-x}y'-e^{-x}y=2xe^{-x}"


"\frac{d}{dx}(ye^{-x})=2xe^{-x}"


"ye^{-x}=\int 2xe^{-x}dx=-2e^{-x}(x+1)+c"


"u(x,y)=c"


"ye^{-x}+2e^{-x}(x+1)=c"


"u(x,y)=ye^{-x}+2e^{-x}(x+1)"


"u(0,y)=y+2=y^2+1"


"y^2-y-1=0"


"y_1=\frac{1-\sqrt{5}}{2}"


"u_1(x,y)=\frac{1-\sqrt{5}}{2}e^{-x}+2e^{-x}(x+1)"


"y_2=\frac{1+\sqrt{5}}{2}"


"u_2(x,y)=\frac{1+\sqrt{5}}{2}e^{-x}+2e^{-x}(x+1)"